OpenStudy (anonymous):

Simplify 5^3√96 (5 cubed times the square root of 96)

OpenStudy (anonymous):

Choices: 20^3√6 10^3√12 7^3√12 2^3√12

OpenStudy (cwrw238):

sqrt96 = sqrt16 * sqrt6 = 4sqrt6

OpenStudy (anonymous):

is it a then?

OpenStudy (anonymous):

it's B

OpenStudy (anonymous):

how did you get to that answer? this section of my course is really confusing me haha

OpenStudy (cwrw238):

5^3 = 125

OpenStudy (anonymous):

okay i got that

OpenStudy (cwrw238):

125 * 4 sqrt6 = 600 sqrt6

OpenStudy (anonymous):

10^3cuberoot 12 = 5*2cuberoot12 =5cuberoot12*2^3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how can 600 sqrt 6 be simplified?

OpenStudy (cwrw238):

yea - I'm looking at that - i'm not at my best today!!!

OpenStudy (anonymous):

haha its okay

OpenStudy (cwrw238):

sqrt600 = sqrt100sqrt6 = 10sqrt6

OpenStudy (anonymous):

can that be simplfied? because 10sqrt6 isnt a choice

OpenStudy (anonymous):

what are your choices?

OpenStudy (anonymous):

Choices: 20^3√6 10^3√12 7^3√12 2^3√12

OpenStudy (anonymous):

what is your question again?

OpenStudy (anonymous):

Simplify 5^3√96

OpenStudy (cwrw238):

i think your interpretation of question may be wrong

OpenStudy (anonymous):

5^3 is 125 not 600 lol

OpenStudy (anonymous):

ooops I thought u said 600 sqrt 6

OpenStudy (cwrw238):

yea i know - but multiplied by 4 thats 600

OpenStudy (cwrw238):

so it comes to 600 sqrt6

OpenStudy (anonymous):

what multiply by 4?

OpenStudy (cwrw238):

oops 500 - my bad maths!!!

OpenStudy (anonymous):

LOL

OpenStudy (cwrw238):

its 500 sqrt6

OpenStudy (anonymous):

Choices: 20^3√6 10^3√12 7^3√12 2^3√12

OpenStudy (anonymous):

yea

OpenStudy (cwrw238):

but thats still not one of the choices!!!

OpenStudy (anonymous):

i know lol

OpenStudy (cwrw238):

this is a comedy of errors

OpenStudy (anonymous):

Hahaha

OpenStudy (cwrw238):

- mostly mine

OpenStudy (anonymous):

me too for not understanding how to do math lol

OpenStudy (anonymous):

let's conclude there's printing error

OpenStudy (cwrw238):

sqrt 96 is approximately 10 so that makes the value about 1250

OpenStudy (anonymous):

1224.74 to be exact

OpenStudy (cwrw238):

right first choice = 8000 sqrt6 - much higher second = 1000 * 3.5 third = 340 * 3.5 = thats close fourth - too low

OpenStudy (cwrw238):

none of the choices match

OpenStudy (anonymous):

Yea, even the third one is not close enough